OpenStudy (prettygirl_shynice):

may someone help with math i need help explaing like how i got answer i know the answer the answer but how i explaini know the answer its b but how i explain 21 to x and 7 to 10 a.24 b.30 c.40

OpenStudy (anonymous):

\[\frac{21}{x}=\frac{7}{10}\]?

OpenStudy (prettygirl_shynice):

okay

OpenStudy (anonymous):

you can do it by thinking since \(7\times 3=21\) then \(10\times 3=30\)

OpenStudy (prettygirl_shynice):

thank u because i was lost

OpenStudy (prettygirl_shynice):

can u explain step by step

OpenStudy (anonymous):

or you can do \[\frac{x}{21}=\frac{10}{7}\iff x=21\times \frac{10}{7}=3\times 10=30\] either way

OpenStudy (prettygirl_shynice):

okay and my answer still is 30

OpenStudy (anonymous):

yes

OpenStudy (prettygirl_shynice):

im srry didnt mean to bother u

OpenStudy (anonymous):

lol no problem

OpenStudy (prettygirl_shynice):

how do u write your fraction like that on the computer

OpenStudy (anonymous):

magic the magic of \(\LaTeX\)

OpenStudy (anonymous):

but you can do it with the equation tool below

OpenStudy (prettygirl_shynice):

okay im about to try

OpenStudy (prettygirl_shynice):

thank u

OpenStudy (anonymous):

yw

OpenStudy (prettygirl_shynice):

lol